Solve the following system of linear equations:
x + 2y - z = 5
2x - y + z = 2
y + z = 2

Found 2 solutions by Alan3354, vic21:
Answer by Alan3354(69443)

x + 2y - z = 5
2x - y + z = 2
y + z = 2
-----------------------
x + 2y - z = 5
2x - y + z = 2
---------------- Add
3x + y = 7 Eqn A
================
x + 2y - z = 5
y + z = 2
------------------ Add
x + 3y = 7 Eqn B
9x + 3y = 21 Eqn A times 3
------------------------------ Subtract
-8x = -14
x = 7/4
-------
x + 3y = 7 Eqn B
3y = 21/4
y = 7/4
------------
y + z = 2
z = 1/4

Answer by vic21(4)

x = 7/4
y = 7/4 too
then z= 1/4
just use the elimination method.